# ACEBench: Who Wins the Match Point in Tool Learning?

### v2

@@ Abstract
Large language models (LLMs) have demonstrated significant potential in decision-making and reasoning, especially when combined with various tools to effectively solve complex problems. However, existing evaluation systems for assessing LLM function calling capabilities have several limitations: (1) limited evaluation scenarios, lacking assessments in real multi-turn dialogue contexts; (2) narrow evaluation dimensions, lacking detailed assessments for fine-grained function calls; (3) relying on LLMs or real API executions for result evaluation, which introduces significant overhead. To address these issues, we propose a comprehensive evaluation system named ACEBench. This system is meticulously designed to encompass a wide spectrum of function calling scenarios. Moreover, it categorizes these scenarios into three primary types according to the evaluation methodology: Normal, Special, and Agent. Normal evaluates function calls in basic scenarios; Special evaluates function calls in scenarios with vague or incomplete instructions; Agent introduces multi-agent interactions to simulate function calling evaluation in real-world multi-turn interactions. We conducted extensive experiments on ACEBench, analyzing various LLMs in-depth and performing a more granular analysis of error causes across different data types.
ACEBench: Who Wins the Match Point in Tool Learning?
@@ 1 Introduction
Large Language Models (LLMs) such as GPT-4 Achiam et al. (2023), have demonstrated exceptional performance across numerous natural language processing tasks Naveed et al. (2023). Tool learning which enables LLMs to tackle more complex problems Qu et al. (2025), has emerged as a promising paradigm for enhancing their ability to handle highly intricate tasks. On one hand, integrating tools into LLMs can enhance capabilities in multiple domains, for example, ToolTransformer Schick et al. (2023) enhances the ability of LLMs to solve complex problems by utilizing tools such as the Wikipedia search engine, calculators, and calendars. On the other hand, adopting a tool learning paradigm can improve response robustness and the transparency of the generation process, thereby increasing explainability and user trust, as well as enhancing system robustness and adaptability. As this field continues to evolve, it is essential to conduct a comprehensive evaluation of all aspects of tool utilization, particularly in complex real-world scenarios.
While several studies have focused on evaluating tool learning Yan et al. (2024); Guo et al. (2024); Wang et al. (2024a); Qin et al. (2023); Wang et al. (2024b), there are still some shortcomings in the existing tool-learning benchmarks. Firstly, existing tool learning benchmarks lack multi-turn dialogue evaluation in real-world scenarios. For example, the multi-turn dialogues in BFCL Yan et al. (2024) and HammerBench  Wang et al. (2024a) are composed of predefined fixed content combinations, which fail to genuinely reflect the interactive process between users and agents in real-life situations. Secondly, current tool learning benchmarks Qin et al. (2023); Guo et al. (2024); Huang et al. (2023); Li et al. (2023) for LLMs lack fine-grained evaluation and personalized data assessment. Therefore, further categorizing data allows us to better understand the impact of different data types on the model’s tool invocation capabilities. Additionally, existing tool learning benchmarks Qin et al. (2023); Guo et al. (2024); Wang et al. (2024b) ignore the assessment of special cases or the evaluation methods are simplistic Yan et al. (2024), as user instructions in real life are not always perfect and may lack necessary parameters, contain error params, or exceed the tool’s capabilities. The model’s ability to recognize and handle these issues is also a crucial aspect of evaluation. Lastly, evaluation costs are high  Qin et al. (2023); Guo et al. (2024), as many studies rely on GPT for evaluations, resulting in significant expenses that hinder large-scale adoption.
Table 1: Comparison of benchmarks across different evaluation criteria. "AutoEval" denotes result evaluation without relying on LLMs. "Imperfect" refers to incomplete or unclear user instructions. "Interaction" refers to the dynamic interaction between the model and the user. "Atom" refers to analyzing from the atomic-level capabilities. "Preference” refers to the inclusion of an individual’s specific likes or choices.
Benchmark AutoEval Imperfect Interaction Atom Preference
@@
To address these deficiencies, we propose a comprehensive tool learning benchmark ACEBench, comprising following categories:
Normal: The normal data consists of fixed question-answer pairs and covers various scenarios, including single-turn single function calls, single-turn parallel function calls, and multi-turn dialogues. It also includes personalized scenarios, such as candidate function lists based on similar APIs and personalized data reflecting user preference. Additionally, normal incorporates fine-grained evaluations, analyzing tool invocation from the perspective of atomic capabilities.
Special: Includes imperfect user instructions, such as instructions lacking necessary parameters, containing incorrectly formatted parameters, or questions exceeding the capabilities of candidate functions.
Agent: Based on many real-world scenarios, abstracted to construct multi-turn, multi-step tool invocation scenarios, divided into multi-turn and multi-step cases depending on whether the user participates in the dialogue process.
The three categories above cover most of the tool usage scenarios for LLMs, and for different evaluation scenarios, we have introduced multiple evaluation metrics (details can be found in Section 3.2). Our main contributions are as follows:
  • Comprehensive Benchmark Evaluation: We propose a comprehensive benchmark for LLMs tool learning, covering various real-life scenarios, including more fine-grained evaluation perspectives and assessments under imperfect instructions, and featuring more stable evaluation metrics.
  • Sandbox Environment and Automated Evaluation System: We develop a sandbox environment construction scheme for multi-turn, multi-step tool invocation based on real-world scenario abstraction and build an end-to-end automated evaluation system.
  • Extensive Experimental Validation: Through extensive experiments, we demonstrate that our benchmark provides a more comprehensive analysis with greater distinction, offering a clearer evaluation of the tool learning capabilities of various models.
@@ 2 Related Works
The emerging trend of leveraging the tool-learning capabilities of LLMs in real-world applications underscores the need for comprehensive evaluations of their performance and effectiveness. Despite recent advancements, existing benchmarks for evaluating the tool learning capabilities of LLMs still exhibit notable limitations.
Stable ToolBench Guo et al. (2024) addresses the issue of unstable external APIs by employing a virtual API server, but its reliance on large models for evaluation results in high costs and scalability challenges. BFCL Yan et al. (2024) introduces a benchmark for tool use in multi-turn dialogue scenarios, yet it assembles dialogues from fixed content, failing to capture the dynamic and adaptive nature of real-world interactions. Similarly, Taubench Alanko et al. (2022) evaluates language agents’ ability to engage with human users while adhering to domain-specific rules, but its narrow focus on just two scenarios limits its generalizability across diverse tasks. HammerBench Wang et al. (2024a) improves upon this by incorporating datasets derived from popular mobile applications and merging dialogues to simulate typical question-answer trajectories. However, like BFCL, its multi-turn dialogues are simplistic concatenations of pre-defined content, which do not reflect the complexities of real-world conversational dynamics. In addition, some benchmarks Qin et al. (2023); Guo et al. (2024) rely on large language models (LLMs) for result evaluation, leading to high costs and unstable operations.
@@ 3.1 Dataset
We construct two versions of the dataset, one in Chinese and one in English, ensuring an equal distribution of data types across both versions.
@@ 3.1.1 Data Generation
Our data creation process is categorized into two types: the creation process for Agent data is detailed in Appendix E, and the creation process for other types of data, which involves the following two steps: API synthesis and dialogue generation.
API Synthesis. Considering the evaluation stability, we use synthetic APIs to construct the evaluation dataset. We use real APIs from various real-world scenarios as reference during synthesis to enhance authenticity. The left part in Figure 1 illustrates the generation of APIs.
To ensure the richness of the generated APIs, we employ a self-evolution approach for API generation Liu et al. (2024b). This method involves constructing a hierarchical API context tree to guide the synthesis process, encompassing a wide range of API domains and functionalities. Initially, we leveraged API-related documents from real world, such as technical manuals, API documentation, product specifications to prompt the LLM to extract relevant domains and their associated functionalities or cases from each document. As the generation process progressed, child nodes were recursively added to the context tree, gradually expanding its structure. By the end of the process, the context tree is thoroughly developed, ensuring both the depth and breadth of the evaluation dataset, and guaranteeing comprehensive coverage of the potential API domains and functionalities. With the API context trees, we maintain a large API pool for subsequent dialog generation.
Dialogue Construction. We use two different dialogue generation pipelines, both built on the constructed API pool from which three to six candidate APIs are selected for each evaluation instance. For most cases, APIs are chosen randomly. However, for instances requiring specific functionality (e.g., similar APIs or multi-turn scenarios), advanced methods including graph-based sampling Wang et al. (2024d) are used. Simple cases or those with predefined functionality use a template-based generation, where a single generator produces dialogues to ensure consistency. For more complex scenarios, we employ a multi-agent dialogue pipeline, where three agents (user, assistant, and tool) role-play to simulate real-world interactions. Both pipelines are supported by carefully hand-crafted examples to ensure comprehensive coverage and diversity. The middle part in Figure 1 illustrates one of dialogue generation pipelines.
Figure 1: Pipeline for Construction of Evaluation Dataset. The module on the far left is the API Synthesis Module. The module in the middle is the Dialogue Generation Module based on the API Pool. The module on the far right is the Dialogue Data Quality Inspection Module.
@@ 3.1.2 Multi-Stage Data Verification
Except for the Agent category, all of our data are directly generated by large language models (LLMs). Despite the fact that the quality of data generated by LLMs has been continuously improving and the number of errors has been decreasing as the capabilities of LLMs are enhanced, due to the prevalent problem of hallucination, data directly generated by LLMs inevitably has certain issues. For instance, the answers may not be in line with the questions, or the answer criteria may be ambiguous. As shown in the right part in Figure 1, we have established a multi-stage data verification scheme.
The quality inspection scheme consists of two stages: the fully automated quality inspection stage and the human quality inspection stage based on LLM feedback. They are respectively used for quickly filtering out obvious error cases and carefully selecting controversial error cases.
Automated Quality Inspection
The data generated by the model must first go through a rule-based quality inspection module, which draws inspiration from the design implemented in ToolACE  Liu et al. (2024b). This module evaluates the data across four crucial dimensions: clarity of API definitions, executability of function calls, accuracy of dialogues, and consistency of data samples, all guided by a meticulously curated set of rules. By applying this rule-based quality inspection, it is possible to filter out the majority of formatting errors, spelling mistakes, and other similar issues, thereby initially enhancing the overall quality of the dataset.
Following the rule-based quality inspection module, the dataset undergoes a model-based quality verification module. This module employs state-of-the-art LLMs to identify more subtle and implicit high-order semantic errors within the data. However, it has been observed that when directly querying these models about the quality of the data (i.e., whether it is good or bad), the responses tend to be ambiguous. To address this issue, we have decomposed the main question into multiple sub-questions, such as "Can the tool selected by the assistant effectively resolve the user’s problem?" and "Are specific constraints and formats of parameters adhered to according to their definitions?". By breaking down the evaluation into these detailed inquiries, we are able to gather more precise feedback on various aspects of the data. Subsequently, these individual assessments are aggregated through a voting mechanism to derive a final verdict on the quality of the data. This comprehensive approach enables us to perform thorough filtering of the dataset, ensuring its reliability and coherence.
Human Quality Inspection Based on LLM Feedback
The initial evaluation module encompasses the utilization of the data set preserved following automatic quality inspection as the test set. Initially, the test set is evaluated by utilizing the three most advanced models currently. These evaluation outcomes will function as auxiliary information to facilitate more effective data screening by human experts. For the test data that is free from any problems, we will retain it directly. In contrast, the data that may be problematic will be marked and added to the error candidate pool, which constitutes around 20% of the total data set in the first turn. This approach ensures a robust filtering mechanism, allowing for iterative refinement and improving the overall quality of the dataset over multiple rounds.
For the entries in the error candidate pool, a two-step human expert review process is initiated. Two experts independently assess each flagged item and provide suggestions for corrections. Subsequently, a third expert consolidates these assessments and executes the necessary corrections, which may include addressing issues related to problem statements, API definitions, and answer. After completing one round of this correction process, the revised test set is re-evaluated using the same models. Following this, a manual verification step is undertaken. Through three rounds of iterative evaluation and correction, the final validated dataset is produced.
@@ 3.1.3 Data Analysis
To demonstrate the breadth and comprehensiveness of ACEBench, we provide a detailed analysis of its test case distributions. This includes an examination of API domains, the number of samples in each test category, and the diversity of API calling patterns.
Domain of APIs. The ACEBench API boasts a comprehensive coverage of 8 major domains and 68 sub-domains, spanning various aspects of daily life, including technology, finance, entertainment, society, healthy, culture, enviroment and others. It offers a rich collection of 4,538 APIs in both Chinese and English. The distribution of these APIs is visualized in the accompanying Figure 2. By exposing the model to such a diverse set of APIs, we can effectively assess its ability to generalize across different API contexts, demonstrating its comprehension of API documentation and its proficiency in executing API calls.
Figure 2: Distribution of APIs in terms of domains.
Data Composition. ACEBench consists of three categories of test samples: Normal, Agent and Special, where each category is divided into several subcategories according to the patterns. The data composition is visualized in Figure 3, demonstrating a comprehensive coverage of tool-calling capabilities, from simple single-turn tool invocations to complex multi-turn interactions involving users and environments. They include scenarios requiring multiple steps and interactions with the environment, as well as cases where tool calls are infeasible. Single-turn data, which forms the majority, underscores the fundamental importance of basic tool-calling capabilities.
@@
Number of turns and arguments. The test data in ACEBench covers a wide range of complexities. Specifically, we statistically analyzed the number of dialogue turns and the number of arguments in the called functions, which are visualized in  Figure4. The results show that the number of dialogue turns ranges from 1 to 8, encompassing most real-world scenarios. The number of arguments in the called functions can be as many as 7, with most functions having 1 to 3 arguments. These samples with varying numbers of turns and arguments further form a test suite that covers a broader range of difficulties, allowing ACEBench to conduct a more granular evaluation of different models.
Figure 4: Distribution of number turns and arguments.
Figure 5: The Overview of Evaluation Process. The left represents "Normal" evaluation (AST parsing for function and parameter checks), the middle illustrates "Special" evaluation (error type matching), and the right shows "Agent" evaluation (final state and intermediate result assessments with user-model interaction).
3.1.4 Data Categories Description
Normal Data
The Normal Data consists of fixed question-answer pairs, where each question corresponds to a correct function call. It is categorized into the following categories: Single Turn, Multi Turn, Similar API, Preference and Atom. Data examples can be found in Appendix A.
Multi Turn: There are multiple rounds of interactions between the user and the assistant. The conversation can be categorized into two types: (switch) The conversation progresses by changing topics. (adjust) The conversation evolves by refining or modifying the original question.
Similar APIs: The candidate APIs exhibit significant similarity, particularly in terms of their focus on the same topic. This similarity presents a challenge for the assistant, requiring it to effectively distinguish between the APIs and accurately select the most appropriate ones.
Special Data
The Special Data refers to situations where the model is unable to resolve the problem posed in the instruction using the candidate functionsWang et al. (2024c). It is categorized into the following categories: Lose param, Error param, and Exceed Function Ability. Data examples can be found in Appendix A.2.
Lose Param: Refers to situations where the key information required for the function call is missing in the query, such as the absence of "required" parameters.
Error Param: Refers to situations where the instruction contains parameters or names that do not meet the required format or constraints, such as matching a specific pattern or being selected from a predefined list, causing the function call to fail.
Exceed Function Ability: Refers to situations where the instruction exceeds the capabilities of the function, meaning none of the candidate functions are able to resolve the issue.
Agent Data
Agent Data refers to situations where completing a task in an environment built based on real-world scenarios typically requires multiple steps. In such cases, an advanced language model is used to simulate the user role, thereby replicating interactions found in real-world scenarios. This approach aims to evaluate the model’s ability to perform in complex interactions. We define a "multi-step" scenario as one in which the user appears only once in the entire conversation, while a "multi-turn" scenario is one where the user appears multiple times throughout the conversation. Data examples can be found in Figure 6.
Table 2: Comprehensive Accuracy Evaluation of Different Models on ACEBench (Chinese and English Combined)
@@ 3.1.4 Data Categories Description
GPT-4o 0.94 0.85 0.77 0.85 0.83 0.88 0.93 0.64 0.86
GPT-4-Turbo 0.94 0.85 0.78 0.86 0.86 0.88 0.87 0.67 0.85
Qwen-Max 0.91 0.81 0.68 0.83 0.83 0.85 0.74 0.65 0.79
DeepSeek-Chat 0.92 0.85 0.77 0.83 0.83 0.87 0.73 0.35 0.75
GPT-4o-mini 0.87 0.76 0.67 0.77 0.78 0.80 0.79 0.34 0.73
Gemini-1.5-Pro 0.85 0.77 0.65 0.80 0.78 0.79 0.79 0.26 0.71
Claude-3-5-Sonnet 0.77 0.73 0.63 0.71 0.72 0.73 0.78 0.40 0.69
Doubao-Pro-32k 0.80 0.56 0.58 0.76 0.66 0.71 0.55 0.26 0.60
@@
Qwen2.5-Coder-32B-Instruct 0.90 0.81 0.71 0.83 0.81 0.84 0.81 0.61 0.80
Qwen2.5-72B-Instruct 0.87 0.80 0.70 0.83 0.81 0.82 0.76 0.46 0.75
Llama-3.1-70B-Instruct 0.83 0.68 0.64 0.79 0.68 0.76 0.38 0.42 0.61
Qwen2.5-7B-Instruct 0.76 0.61 0.59 0.72 0.67 0.70 0.47 0.14 0.55
DeepSeek-Coder-V2-Lite-Instruct 0.76 0.58 0.47 0.72 0.65 0.67 0.40 0.02 0.50
Qwen2.5-Coder-7B-Instruct 0.76 0.64 0.58 0.74 0.68 0.71 0.22 0.15 0.49
Watt-Tool-8B 0.86 0.70 0.56 0.79 0.64 0.76 0.06 0.03 0.46
Hammer2.1-7b 0.74 0.58 0.40 0.62 0.55 0.63 0.15 0.17 0.43
Llama-3.1-8B-Instruct 0.52 0.40 0.28 0.66 0.46 0.47 0.21 0.06 0.34
Phi-3-mini-128k-Instruct 0.57 0.40 0.23 0.58 0.32 0.47 0.19 0.01 0.33
xLAM-7b-r 0.44 0.22 0.19 0.61 0.00 0.34 0.03 0.09 0.22
Llama-3.2-3B-Instruct 0.39 0.16 0.09 0.42 0.32 0.30 0.10 0.00 0.20
Hammer2.1-3b 0.23 0.12 0.04 0.40 0.20 0.19 0.01 0.01 0.11

### v3

# ACEBench: Who Wins the Match Point in Tool Usage?

@@ Abstract
Large Language Models (LLMs) have demonstrated significant potential in decision-making and reasoning, particularly when integrated with various tools to effectively solve complex problems. However, existing benchmarks for evaluating LLMs’ tool usage face several limitations: (1) limited evaluation scenarios, often lacking assessments in real multi-turn dialogue contexts; (2) narrow evaluation dimensions, with insufficient detailed assessments of how LLMs use tools; and (3) reliance on LLMs or real API executions for evaluation, which introduces significant overhead. To address these challenges, we introduce ACEBench, a comprehensive benchmark for assessing tool usage in LLMs. ACEBench categorizes data into three primary types based on evaluation methodology: Normal, Special, and Agent. "Normal" evaluates tool usage in basic scenarios; "Special" evaluates tool usage in situations with ambiguous or incomplete instructions; "Agent" evaluates tool usage through multi-agent interactions to simulate real-world, multi-turn dialogues. We conducted extensive experiments using ACEBench, analyzing various LLMs in-depth and providing a more granular examination of error causes across different data types.
ACEBench: Who Wins the Match Point in Tool Usage?
@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 Achiam et al. (2023), have demonstrated exceptional performance across numerous natural language processing tasks Naveed et al. (2023); Qu et al. (2025); Mialon et al. (2023). Studies have shown that incorporating tools can significantly expand LLM capabilities, particularly in specialized domains such as mathematics Das et al. (2024); Bulusu et al. (2024); Gou et al. (2023); Veerendranath et al. (2024), programming Xu et al. (2024), and reasoning Chen et al. (2022); Shao et al. (2022); Surís et al. (2023); Yang et al. (2023). On one hand, integrating tools into LLMs can enhance capabilities in multiple domains, for example, ToolTransformer Schick et al. (2023) enhances the ability of LLMs to solve complex problems by utilizing tools. On the other hand, adopting a tool usage paradigm can improve response robustness and the transparency of the generation, thereby increasing explainability and user trust Schick et al. (2023), as well as improving the system’s adaptability. As this field continues to evolve, it is essential to comprehensively evaluate all aspects of tool usage, particularly in complex scenarios.
Table 1: Comparison of benchmarks across different evaluation criteria. "LLM-Free" refers to result evaluation without relying on LLMs. "Robustness" refers to incomplete or unclear user instructions. "Interactiveness" refers to the dynamic interaction between the model and the environment. "Atomic-Level" refers to analyzing from the atomic-level capabilities. "Personalization” refers to the inclusion of personal likes.
Benchmark LLM-Free Robustness Interactiveness Atomic-Level Personalization
@@
While several studies have focused on evaluating tool usage Yan et al. (2024); Guo et al. (2024); Wang et al. (2024a); Qin et al. (2023); Wang et al. (2024b); Zhuang et al. (2023); Lu et al. (2024), there are still some shortcomings in the existing tool-use benchmarks. Firstly, existing benchmarks lack multi-turn dialogue evaluation in real-world scenarios. For example, the multi-turn dialogues in BFCL Yan et al. (2024) and HammerBench  Wang et al. (2024a) are composed of predefined fixed content combinations. Secondly, current tool-use benchmarks Qin et al. (2023); Guo et al. (2024); Huang et al. (2023); Li et al. (2023) lack fine-grained evaluation and personalized data assessment. Additionally, existing benchmarks Qin et al. (2023); Guo et al. (2024); Wang et al. (2024b) ignore the assessment of special cases, or the evaluation methods are simplistic Yan et al. (2024), as user instructions in real life are not always perfectWang et al. (2024c). The model’s ability to recognize and handle these issues is also crucial for evaluation. Lastly, evaluation costs are high  Qin et al. (2023); Guo et al. (2024), as many studies rely on advanced large models for evaluation.
To address these shortcomings, we propose ACEBench, a comprehensive tool-use benchmark that includes the following categories:
Normal. Consists of fixed question-answer pairs and encompasses a variety of scenarios, including single-turn dialogues, multi-turn dialogues, and personalized scenario data. It also includes evaluations of atomic-level capabilities.
Special. Includes imperfect instructions, such as instructions containing incomplete parameters, incorrectly formatted parameters, or questions that are irrelevant to the capabilities of the candidate functions.
Agent. Encompasses real-world scenarios, abstracted to construct multi-turn, multi-step tool invocation scenarios, divided into multi-turn and multi-step cases depending on whether the user participates in the dialogue process.
The three categories above cover most of the tool usage scenarios for LLMs, and detailed explanations of each category can be found in Appendix A. Our main contributions are as follows:
  • Comprehensive Benchmark Evaluation. We propose a comprehensive benchmark for evaluating LLMs’ tool usage, covering various scenarios, including more fine-grained evaluation perspectives and assessments under imperfect instructions, and providing more stable evaluation metrics.
  • Sandbox Environment and Automated Evaluation System. We develop a sandbox environment construction scheme for multi-turn, multi-step tool invocation based on real-world scenario abstraction and build an end-to-end automated evaluation system.
  • Extensive Experimental Validation. Through extensive experiments, we demonstrate our benchmark provides a more comprehensive analysis with greater distinction, offering a clearer evaluation of LLMs’ tool usage.
Figure 1: Evaluation Dataset Construction Pipeline: API Synthesis Module (left), Dialogue Generation Module (middle), Quality Inspection Module (right).
@@ 2 Related Works
The emerging trend of leveraging LLMs’ tool-use capabilities in real-world applications underscores the need for comprehensive evaluations of their performance and effectiveness. Despite recent advancements, existing benchmarks for evaluating the tool-use capabilities of LLMs still have significant limitations
Stable ToolBench Guo et al. (2024) addresses the issue of unstable external APIs by employing a virtual API server, but its dependence on large models for evaluation results in high costs and scalability challenges. BFCL Yan et al. (2024) introduces a benchmark for tool use in multi-turn dialogue scenarios. Yet, it assembles dialogues from fixed content, failing to capture real-world interactions’ dynamic and adaptive nature. Similarly, Taubench Alanko et al. (2022) evaluates language agents’ ability to engage with human users while adhering to domain-specific rules. Still, its narrow focus on just two scenarios limits its generalizability across diverse tasks. HammerBench Wang et al. (2024a) improves upon this by incorporating datasets derived from popular mobile applications and merging dialogues to simulate typical question-answer trajectories. However, like BFCL, its multi-turn dialogues are simplistic concatenations of pre-defined content, which do not reflect the complexities of real-world conversational dynamics. In addition, some benchmarks Qin et al. (2023); Guo et al. (2024) rely on large language models (LLMs) for result evaluation, leading to high costs and unstable operations.
@@ 3.1 Dataset
We construct two versions of the dataset, one in Chinese and the other in English, ensuring an equal distribution of data types across both versions.
@@ 3.1.1 Data Generation
The data construction process is divided into two categories: (1) the creation of Agent data, as detailed in Appendix B.1, and (2) the creation of other types of data, which includes two primary steps: API synthesis and dialogue generation.
API Synthesis. We use real APIs from various real-world scenarios as reference during synthesis to enhance authenticity. To ensure the stability of the data, we use synthetic APIs to construct the evaluation dataset, referencing real-world APIs as a guide. We employ a self-evolution approach by building a hierarchical API context tree to ensure the generated APIs cover a wide range of domains and functionalities Liu et al. (2024b). Initially, we extract relevant information from technical documents to guide the API generation. As the process progresses, the context tree is gradually expanded, ultimately ensuring the depth and breadth of the generated APIs. The left part in Figure 1 illustrates the generation of APIs.
Dialogue Construction. As shown in the middle part in Figure 1, we use two different dialogue generation pipelines built on the constructed API pool from which three to six candidate APIs are selected for each evaluation instance. For most cases, APIs are chosen randomly. However, for instances requiring specific functionality (e.g., similar APIs or multi-turn scenarios), advanced methods, including graph-based sampling Wang et al. (2024d), are used. Simple cases or those with predefined functionality use a template-based generation, where a single generator produces dialogues to ensure consistency. We employ a multi-agent dialogue pipeline for more complex scenarios, where three agents (user, assistant, and tool) role-play to simulate real-world interactions. Both pipelines are supported by carefully hand-crafted examples to ensure comprehensive coverage and diversity. A detailed description for special data construction is provided in Appendix B.2.
@@ 3.1.2 Multi-Stage Data Verification
To address issues like mismatched answers or ambiguous criteria, we have implemented a multi-stage verification process, shown on the right part of Figure 1.
Automated Quality Inspection. The data first undergoes a rule-based quality inspection module, which evaluates four dimensions: clarity of API definitions, executability of function calls, accuracy of dialogues, and consistency of data samples, effectively filtering out formatting and spelling errors. Next, the data enters the model-based quality verification module, which uses LLMs to detect semantic errors, employing a voting mechanism to ensure consistency in evaluation.
Human Quality Inspection. In the initial evaluation, the dataset remaining after automated quality inspection is assessed by three LLMs to assist human experts in data screening. Valid data is retained, while potentially problematic data is placed in the error candidate pool (approximately 20% of the dataset). These flagged entries undergo a two-step expert review process, where two experts independently assess and suggest modifications, and a third expert consolidates feedback, revising problem statements, API definitions, and answers. The revised data is re-evaluated and manually verified, and three rounds of optimization are performed to ensure a high-quality dataset.
@@ 3.1.3 Data Analysis
To demonstrate the breadth and comprehensiveness of ACEBench, we provide a detailed analysis of its test case distributions. This includes an examination of API domains, the number of samples in each category, and the diversity of API patterns.
Figure 2: Distribution of APIs in terms of domains (Top 3 Subcategories for Each Category).
Domain of APIs. The ACEBench API boasts a comprehensive coverage of 8 major domains and 68 sub-domains, spanning various aspects of daily life, including technology, finance, entertainment, society, health, culture, environment, and others. It offers a rich collection of 4,538 APIs in both Chinese and English. The distribution of these APIs is visualized in the accompanying Figure 2.
@@
Data Composition. ACEBench consists of three categories of test samples: Normal, Agent, and Special, where each category is divided into several subcategories. The data composition is visualized in Figure 3, demonstrating a comprehensive coverage of tool-use capabilities, from simple single-turn tool invocations to complex multi-turn interactions involving users and environments. They include scenarios requiring multiple steps and interactions with the environment, as well as cases where tool calls are infeasible.
Figure 4: Distribution of Dialogue Turns and API Argument Numbers.
Number of turns and arguments. The test data in ACEBench covers a wide range of complexities. Specifically, we statistically analyzed the number of dialogue turns and the number of arguments in the called apis, which are visualized in Figure 4. The results show that the number of dialogue turns ranges from 1 to 8, encompassing most real-world scenarios. These samples with varying numbers of turns and arguments further form a test suite that covers a broader range of difficulties, allowing ACEBench to conduct a more granular evaluation of different models.
Figure 5: Overview of Evaluation Process: The left represents ’Normal’ evaluation, which involves AST parsing for function and parameter checks. The middle illustrates ’Special’ evaluation, focused on error type matching. The right shows ’Agent’ evaluation, which assesses final state and intermediate results through user-model interaction.
Table 2: Comprehensive Evaluation of Different Models on ACEBench for Chinese and English Combined (%).
@@
GPT-4o 93.4 84.5 77.0 85.0 83.0 87.6 93.0 63.8 85.4
GPT-4-Turbo 93.2 84.8 77.5 86.0 86.0 88.0 86.7 67.5 84.5
Qwen-Max 91.2 80.5 68.0 83.0 83.0 84.2 74.0 64.3 78.4
GPT-4o-Mini 86.5 76.0 66.5 77.0 78.0 79.9 79.0 33.3 72.5
Gemini-1.5-Pro 84.5 76.8 64.5 80.0 78.0 79.0 78.7 25.5 70.7
Claude-3-5-Sonnet 76.9 72.5 62.5 71.0 72.0 72.9 77.4 39.5 68.9
Doubao-Pro-32k 79.8 55.5 58.0 76.0 66.0 70.7 55.0 25.0 59.4
@@
Qwen2.5-Coder-32B-Instruct 90.2 81.0 71.0 83.0 81.0 84.1 80.7 60.8 79.6
DeepSeek-V3 91.5 84.0 77.0 83.0 83.0 86.5 73.0 34.5 74.8
Qwen2.5-72B-Instruct 86.8 80.3 69.5 83.0 81.0 82.1 75.7 45.0 74.7
Llama-3.1-70B-Instruct 82.5 68.3 63.5 79.0 68.0 75.5 38.3 42.3 60.4
Qwen2.5-7B-Instruct 76.0 60.3 58.5 72.0 67.0 69.4 47.0 13.8 54.8
DeepSeek-Coder-V2-Lite-Instruct 75.2 57.8 46.5 72.0 65.0 66.4 40.3 2.0 49.5
Qwen2.5-Coder-7B-Instruct 76.0 63.8 57.5 74.0 68.0 70.1 22.3 15.5 48.9
Watt-Tool-8B 85.7 69.3 55.5 79.0 64.0 75.6 6.0 2.8 45.7
Hammer2.1-7B 73.7 57.5 40.0 62.0 55.0 62.8 14.7 16.8 42.9
Llama-3.1-8B-Instruct 51.9 39.8 28.0 66.0 46.0 46.6 21.0 5.3 33.4
Phi-3-Mini-128k-Instruct 57.2 39.3 23.0 58.0 32.0 46.5 18.7 0.8 32.0
xLAM-7B-r 43.5 22.0 19.0 61.0 0.0 33.7 2.7 8.8 21.6
Llama-3.2-3B-Instruct 38.7 15.3 9.0 42.0 32.0 29.6 9.4 0.0 19.6
Hammer2.1-3B 22.4 11.5 3.5 40.0 20.0 18.7 1.0 1.5 11.3